a no. consist of two digits of which ten's diit exceed the unit's digit by 6 . the no. itself is equal to the ten times the sum of its digits. find the number.
Answer by checkley71(8403)

LET Y (UNITS DIGIT) PLUS 6 EQUAL THE TENS DIGIT X
Y=X-6
10(X+Y)=10X+Y REPLACE Y WITH (X-6)
10(X+X-6)=10X+(X-6)
10(2X-6)=10X+X-6
20X-60=11X-6
20X-11X=60-6
9X=54
X=54/9
X=6 WHICH IS THE TENS DIGIT
Y=6-6=0 FOR THE UNITS DIGIT'
PROOF
10(6+0)=10*6+0
60=60